Chapter 2, Practice Problem 2/036 Water is ejected from the water nozzle with a speed Vo = 49.5 ft/sec. For what value of theWater is ejected from the water nozzle with a speed v0 = 49.5 ft/sec. For what value of the angle θ will the water land closest to the wall after clearing the top? Neglect the effects of wall thickness and air resistance. What distance x to the right of point B does the water land?
Answer #1

Imagine water stream as particles, its location in (x,y) planeis, (x,y)=( 69,4.9) Along x axis , speed of water is uniform

(2) Using 8 = 54.56, time taken by water to reach the land is, 0 =1.0 +(49.5t ) sin 8 - 0 =1.0+(49.5t ) sin 54.56* - }(32.2